OpenStudy (anonymous):

Sketch a function with f'(x)>0 for x<2, f'(2)=0 and f'(x)<0 when x>2

OpenStudy (campbell_st):

so there is a stationary point at x = 2 the slope of the tangent to the curve is positive for x < 2 and the slope of the tangent is negative for x > 2 I hope this helps...

OpenStudy (anonymous):

Yes I know that but still having troubles with it, please help?

OpenStudy (campbell_st):

ok... well to me, based on the information, its a concave down parabola with a vertex at x = 2 any you have no information about f(2)

OpenStudy (campbell_st):

I hope this helps you... the curve may be negative definite, touch the axis at x = 2 or have 2 unequal real roots

OpenStudy (anonymous):

Could you please tell me how you know the vertex is 2, and that it is a concave down?

OpenStudy (campbell_st):

well the 1st derivative.... stationary points in any curve occur when the 1st derivative is equal to zero and solved... e.g. f(x) = x^2 + 6x f'(x) = 2x + 6 let f'(x) = 0 then 2x + 6 = 0 so x = -2 is a stationary point... or has a slope of zero at x = -3 so looking at the slope of the tangents in your question 

OpenStudy (anonymous):

Ohhh ok I see it now

OpenStudy (anonymous):

Thank You for your time